At what distance must two equal sphere be place so that the force of attraction between them would be 3.35×10-6N if the mass of the sphere is 90kg each

1
Expert's answer
2022-01-14T09:45:21-0500

By the law of universal gravitation:


"F=G\dfrac {m_1m_2}{R^2}→R=\sqrt{\dfrac{Gm_1m_2}{F}}=40\text{ cm}."